OpenStudy (anonymous):

A circle has a diameter that starts at (3,7) and ends at (7,3) . What is the radius of circle rounded to nearest tenth

OpenStudy (gorv):

diameter = distance between these 2 points \[=\sqrt{(7-3)^{2}+(3-7)^2}=\sqrt{16+16}=\sqrt{2*16}=4*\sqrt{2}\]

OpenStudy (gorv):

radius=diameter/2=4 (sqrt2)/2=2*sqrt2